OpenStudy (anonymous):

Find the slope of the two points: (15, 35) and (195, 50) & simplify if possible?

OpenStudy (anonymous):

slope = (50-35)/ (195-15) = 1/12

OpenStudy (anonymous):

You're cool.

OpenStudy (anonymous):

ROFLMFAO

OpenStudy (anonymous):

Some people and their children

OpenStudy (anonymous):

the formula for problems like this are (y2-y1)/(x2-x1)

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

That should almost always give you the slop

OpenStudy (anonymous):

anytime